An astronomical telescope has a length of 44 cm and tenfold magnification. The focal length of the objective lens is


        (a)    4 cm


        (b)    40 cm


        (c)    44 cm


        (d)    440 cm


Grade:upto college level

1 Answers

Aditi Chauhan
10 years ago

(b)

Given: Length of astronomical telescope (f0 + fe) = 44 cm and magnification

 (f_0/f_e )=10.

From the given magnification, we find that f0 = 10 fe.

Therefore, 10 fe + fe = 44 or 11 fe = 44 or fe = 4.

And focal length of the objective (f0) = 44 - fe = 44 - 4 = 40 cm.